OpenStudy (anonymous):

Please help! What are the x-coordinates of the solutions to this system of equations? x2 + y2 = 16 y = x + 4

OpenStudy (anonymous):

put y=x+4 in eq 1 x2+(x+4)^2=16 =>x2+x2+16+4x=16 =>2x2+4x=0 =>2x(x+4)=0 =>x=0,-4

OpenStudy (anonymous):

Thank you so much! I have a question (x+4)(x+4)= x^2+8x+16 right? So how I get how you got x^2+x^2=2x^2 but what about the 8x

OpenStudy (anonymous):

(x+4)(x+4)=(x+4)^2 so use it (a+b)^2=a2+b2+2ab and sorey...you are right.its 8x not 4x.. so factrs regarding x=0,-8

OpenStudy (anonymous):

my mistake..

OpenStudy (anonymous):

Ohhh..thank you so much!